Solve 𝑧𝑝 = βˆ’π‘₯.

1
Expert's answer
2022-04-06T11:57:23-0400

"\ud835\udc67\ud835\udc5d = \u2212\ud835\udc65"

"p=\frac{dz}{dx}"

"𝑧\frac{dz}{dx} = −𝑥"

"zdz=-xdx"

"\frac{z^2}{2}=-\frac{x^2}{2}+\frac C2"

"z^2+x^2=C"

Answer: "z^2+x^2=C" .